OpenStudy (anonymous):

Can some one explain me this step in adding and subtracting? √16.2 + √4.2 √16 * √2 + √4 * √2 4√2 + 2√2 -----> What happen in here?

jhonyy9 (jhonyy9):

second : 4sqrt2 +2sqrt2 = 6sqrt2 4sqrt2 +2sqrt2 ----- 6sqrt2

OpenStudy (anonymous):

Thanks jhonyy9 I have the answer I just don't know how to explain the process in words in that part.

OpenStudy (anonymous):

\[4\sqrt2+2\sqrt2=\sqrt2(4+2)=\sqrt2\cdot6=6\sqrt2\] do you understand now?

jimthompson5910 (jim_thompson5910):

In that step, you took the square root of 16 to get 4 (in the terms before the plus) and the square root of 4 to get 2 (in the terms after the plus)

OpenStudy (anonymous):

Thanks ...I know is kind of easy for your guys but for me I feel my brain is turning into scramble eggs. :)